OpenStudy (anonymous):

Which table of values goes with the equation y = 3x2 + 1? x y -2 -11 -1 -2 0 1 1 4 2 13 x y -2 13 -1 4 0 1 1 4 2 13 x y -2 13 -1 4 0 1 1 -2 2 -11 x y -2 -11 -1 -2 0 1 1 -4 2 13

OpenStudy (klimenkov):

What is the problem for you?

OpenStudy (anonymous):

huh ?

mathslover (mathslover):

can u please edit the question and remove spaces?

mathslover (mathslover):

y = 3x2 + 1 x = \(\large{\sqrt{\frac{y-1}{3}}}\) if y = 0 then : x = \(\large{\sqrt{\frac{-1}{3}}}\)

mathslover (mathslover):

I didn't get the question correctly .. i think